Determine the molar solubility of Mg(OH)2 in a solution buffered at a pH of 7.
1
Expert's answer
2017-07-28T15:45:07-0400
Ksp(Mg(OH)2) = [Mg+2] * [OH-]^2 = 1.8*10^(-11).
pH = 7; pOH = -lg[OH-] =14 - pH = 7
[OH-] = 10^(-7) mol/l

Mg(OH)2 = Mg+2 + 2OH- ==> 1 mol (Mg(OH)2) = 1 mol (Mg+2)

[Mg+2] = Ksp(Mg(OH)2) / [OH-]^2 = 1.8*10^(-11) / (10^(-7))^2 = 1.8*10^(-11) / 10^(-14) = 1.8*10^3 (mol/l)